# SeaFormer++: Squeeze-enhanced Axial Transformer for Mobile Visual Recognition

### v5

@@ 1 Introduction
As a fundamental problem in computer vision, semantic segmentation aims to assign a semantic class label to each pixel in an image. Conventional methods rely on stacking local convolution kernel [1] to perceive the long-range structure information of the image.
@@
Since the introduction of Vision Transformers [2], the landscape of semantic segmentation has significantly revolutionized. Transformer-based approaches [3, 4] have remarkably demonstrated the capability of global context modeling. However, the computational cost and memory requirement of Transformer render these methods unsuitable on mobile devices, especially for high-resolution imagery inputs.
Following conventional wisdom of efficient operation, local/window-based attention [5, 6, 7, 8], Axial attention [9, 10, 11], dynamic graph message passing [12, 13] and some lightweight attention mechanisms [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] are introduced.
@@
Figure 1: Left: Latency comparison with Transformer [27], MixFormer [28], ACmix [29], Axial attention [10] and local attention [5]. It is measured with a single module of channel dimension 64 on a Qualcomm Snapdragon 865 processor. Right: The mIoU versus latency on the ADE20K val set. MV2 means MobileNetV2 [30]. MV3-L means MobileNetV3-Large [31]. MV3-Lr denotes MobileNetV3-Large-reduce [31]. The latency is measured on a single Qualcomm Snapdragon 865, and only an ARM CPU core is used for speed testing. No other means of acceleration, e.g., GPU or quantification, is used. For figure Right, the input size is 512×512. SeaFormer achieves superior trade-off between mIoU and latency.
@@
However, these advances are still insufficient to satisfy the design requirements and constraints for mobile devices due to the high latency on the high-resolution inputs (see Figure 1). Recently there is a surge of interest in building a Transformer-based semantic segmentation. In order to reduce the computation cost at high resolution, TopFormer [32] dedicates to applying the global attention at a $1/64$ scale of the original input, which definitely harms the segmentation performance.
To solve the dilemma of high-resolution computation for pixel-wise segmentation task and low latency requirement on the mobile device in a performance harmless way, we propose a family of mobile-friendly Transformer-based semantic segmentation model, dubbed as squeeze-enhanced Axial Transformer (SeaFormer), which reduces the computational complexity of axial attention from $\mathcal{O}((H+W)HW)$ to $\mathcal{O}(HW)$, to achieve superior accuracy-efficiency trade-off on mobile devices and fill the vacancy of mobile-friendly efficient Transformer.
@@
The core building block squeeze-enhanced Axial attention (SEA attention) seeks to squeeze (pool) the input feature maps along the horizontal/vertical axis into a compact column/row and computes self-attention. We concatenate query, keys and values to compensate the detail information sacrificed during squeeze and then feed it into a depth-wise convolution layer to enhance local details.
@@
Coupled with a light segmentation head, our design (see Figure 2) with the proposed SeaFormer layer in the small-scale feature is capable of conducting high-resolution image semantic segmentation with low latency on the mobile device. As shown in Figure 1, the proposed SeaFormer outperforms other efficient neural networks on the ADE20K dataset with lower latency. In particular, SeaFormer-Base is superior to the lightweight CNN counterpart MobileNetV3 (41.0 vs.33.1 mIoU) with lower latency (106ms vs.126ms) on an ARM-based mobile device.
@@
A preliminary version of this work was presented in ICLR 2023 [33]. In this paper, we have further extended our conference version as follows: (i) We propose an adaptive squeeze and expand method in Squeeze-axial attention, using a learnable mask to map all tokens of query/key/values to a single token in each row and column; (ii) We present a feature upsampling based multi-resolution distillation approach, further reducing the inference latency of the proposed framework.
@@ 2.1 Combination of Transformers and convolution
Convolution is relatively efficient but not suitable to capture long-range dependencies and vision Transformer has the powerful capability with a global receptive field but lacks efficiency due to the computation of self-attention. In order to make full use of both of their advantages, MobileViT [26], TopFormer [32], LVT [34], Mobile-Former [35], EdgeViTs [36], MobileViTv2 [37], EdgeFormer [38] and EfficientFormer [39] are constructed as efficient ViTs by combining convolution with Transformers. MobileViT, Mobile-Former, TopFormer and EfficientFormer are restricted by Transformer blocks and have to trade off between efficiency and performance in model design. LVT, MobileViTv2 and EdgeViTs keep the model size small at the cost of relatively high computation, which also means high latency.
@@ 2.2 Axial attention and variants
Axial attention [9, 10, 11] is designed to reduce the computational complexity of original global self-attention [27]. It computes self-attention over a single axis at a time and stacks a horizontal and a vertical axial attention module to obtain the global receptive field. Strip pooling [14] and Coordinate attention [40] uses a band shape pooling window to pool along either the horizontal or the vertical dimension to gather long-range context. Kronecker Attention Networks [41] uses the juxtaposition of horizontal and lateral average matrices to average the input matrices and performs attention operation. These methods and other similar implementations provide performance gains partly at considerably low computational cost compared with Axial attention. However, they ignore the lack of local details brought by the pooling/average operation.
@@ 2.3 Mobile semantic segmentation
The mainstream of efficient segmentation methods are based on lightweight CNNs. DFANet [42] adopts a lightweight backbone to reduce computation cost and adds a feature aggregation module to refine high-level and low-level features. ICNet [43] designs an image cascade network to speed up the algorithm, while BiSeNet [44, 45] proposes two-stream paths for low-level details and high-level context information, separately. Fast-SCNN [46] shares the computational cost of the multi-branch network to yield a run-time fast segmentation CNN. TopFormer [32] presents a new architecture with a combination of CNNs and ViT and achieves a good trade-off between accuracy and computational cost for mobile semantic segmentation. However, it is still restricted by the heavy computation load of global self-attention.
@@ 3.1 Overall architecture
Inspired by the two-branch architectures [45, 46, 47, 48, 35], we design a squeeze-enhanced Axial Transformer (SeaFormer) framework. As is shown in Figure 2, SeaFormer consists of these parts: shared STEM, context branch, spatial branch, fusion block and light segmentation head. For a fair comparison, we follow TopFormer [32] to design the STEM. It consists of one regular convolution with stride of 2 followed by four MobileNet blocks where stride of the first and third block is 2. The context branch and the spatial branch share the produced feature map, which allows us to build a fast semantic segmentation model.
@@ 3.2 Squeeze-enhanced Axial attention
To improve the efficiency, there are some works [6, 9, 10] computing self-attention within the local region. We show two most representative efficient Transformer in Equation 23. Equation 2 is represented by window-based attention [5] successfully reducing the time complexity to $\mathcal{O}(m^{2}HW(C_{qk}+C_{v}))=\mathcal{O}(HW)$, where $\mathcal{N}_{m\times m}(o)$ means the neighbour $m\times m$ positions of $o$, but loosing global receptiveness. The Equation 3 is represented by Axial attention [10], which only reduces the time complexity to $\mathcal{O}((H+W)HW(C_{qk}+C_{v}))=\mathcal{O}((HW)^{1.5})$, where $\mathcal{N}_{H\times 1}(o)$ means all the positions of the column of $o$; $\mathcal{N}_{1\times W}(o)$ means all the positions of the row of $o$.
@@ 3.2.5 Architecture details and variants
Figure 5: The overall pipeline of multi-resolution distillation based on feature up-sampling. Fusion block and light segmentation head are consistent with those in Figure 2. MV2(E=4) denotes MobileNetV2 block with expansion ratio of 4, and the default kernel size for depth-wise convolution is 5.
@@
Table 1: Architectures for semantic segmentation. [Conv, 3 ,16, 2] denotes regular convolution layer with kernel of 3, output channel of 16 and stride of 2. [MB, 3, 4, 16, 2] means MobileNetV2 [30] block with kernel of 3, expansion ratio of 4, output channel of 16 and stride of 2. [Sea, 2, 4] refers to SeaFormer layers with number of layers of 2 and heads of 4.
@@ 3.3 Multi-resolution distillation based on feature up-sampling
In dealing with dense prediction tasks, accurately extracting semantic information and spatial details from images often necessitates the input of high-resolution images. This approach undoubtedly increases the computational burden on the model, especially in resource-constrained environments such as mobile devices or real-time applications, where the demand for high computation becomes even more significant. In recent years, to alleviate this issue, some research [49, 50] has proposed methods utilizing multi-scale distillation techniques. In these methods, a teacher model that takes high-resolution images as input is used to guide a student model that takes low-resolution images as input to learn. This strategy allows the student model to produce reasonable results even with low-resolution inputs, thereby significantly reducing computational costs and increasing inference speed.
@@
Furthermore, the upscaled features from the student model are not only used for alignment with the teacher model’s features but are also integrated into the overall semantic segmentation framework, serving as spatial branch features in the feature fusion process. The specific method of this feature fusion and the involved modules will be detailed in Section 3.1. Through this series of steps, the fused features are finally fed into a lightweight segmentation head to complete the semantic segmentation task.
@@ 3.4.2 Loss function
Table 2: Results of semantic segmentation on ADE20K val set, * indicates training batch size is 32. KD means knowledge distillation. The latency is measured on a single Qualcomm Snapdragon 865 with input size 512×512, and only an ARM CPU core is used for speed testing. MobileNetV3-Lr means MobileNetV3-Large-reduce. HRNet-W18S means HRNet-W18-Small. References: MobileNetV2 [30], MobileNetV3 [31], HRNet [51], TopFormer [32], ConvMLP [52], Semantic FPN [53], EfficientNet [54], DeepLabV3+ and Lite-ASPP [55], SegFormer [4], ResNet [56], ShuffleNetV2-1.5x [57].
@@ 4 Experiments
We evaluate our method on semantic segmentation, image classification and object detection tasks. First, we describe implementation details and compare results with state of the art. We then conduct a series of ablation studies to validate the design of SeaFormer. Each proposed component and important hyper-parameters are examined thoroughly.
@@ 4.1.1 Dataset
We perform semantic segmentation experiments over ADE20K [58], CityScapes [59], Pascal Context [60] and COCO-Stuff [61]. The mean of intersection over union (mIoU) is set as the evaluation metric. We convert full-precision models to TNN [62] and measure latency on an ARM-based device with a single Qualcomm Snapdragon 865 processor.
@@ 4.1.2 Implementation details
We set ImageNet-1K [63] pretrained network as the backbone, and training details of ImageNet-1K are illustrated in the last subsection. For semantic segmentation, the standard BatchNorm [64] layer is replaced by synchronized BatchNorm. Our implementation is based on public codebase mmsegmentation [65]. We follow the batch size, training iteration scheduler and data augmentation strategy of TopFormer [32] for a fair comparison.
@@
The initial learning rate is 0.0005 for batch size 32 or 0.00025 for batch size 16. The weight decay is 0.01. A “poly” learning rate scheduled with factor 1.0 is adopted.
@@ 4.2 Comparison with state of the art
Table 2 shows the results of SeaFormer and previous efficient backbones on ADE20K val set. The comparison covers Params, FLOPs, Latency and mIoU. As shown in Table 2, SeaFormer outperforms these efficient approaches with comparable or less FLOPs and lower latency. Compared with specially designed mobile backbone, TopFormer, which sets global self-attention as its semantics extractor, SeaFormer achieves higher segmentation accuracy with lower latency. And the performance of SeaFormer-B++ surpasses MobileNetV3 by a large margin of +8.3% mIoU with lower latency (-16%). The results demonstrate that our SeaFormer layers improve the representation capability significantly.
@@
We compare SeaFormer with the previous approaches on COCO-Stuff validation set in Table 5. From the results, it can be seen that SeaFormer-S++ is +1.2% mIoU higher (34.04 vs.32.82) than SeaFormer-S with similar computation cost.

### v6

@@ 1 Introduction
As a fundamental problem in computer vision, semantic segmentation aims to assign a semantic class label to each pixel in an image. Conventional methods rely on stacking local convolution kernel Long et al. (2015) to perceive the long-range structure information of the image. Since the introduction of Vision Transformers Dosovitskiy et al. (2021), the landscape of semantic segmentation has significantly revolutionized. Transformer-based approaches Zheng et al. (2021); Xie et al. (2021) have remarkably demonstrated the capability of global context modeling. However, the computational cost and memory requirement of Transformer render these methods unsuitable on mobile devices, especially for high-resolution imagery inputs. Following conventional wisdom of efficient operation, local/window-based attention Luong (2015); Liu et al. (2021); Huang et al. (2021a); Yuan et al. (2021), Axial attention Huang et al. (2019b); Ho et al. (2019); Wang et al. (2020a), dynamic graph message passing Zhang et al. (2020, 2023) and some lightweight attention mechanisms Hou et al. (2020); Li et al. (2021a, b, 2020); Liu et al. (2018); Shen et al. (2021); Xu et al. (2021); Cao et al. (2019); Woo et al. (2018); Wang et al. (2020b); Choromanski et al. (2021); Chen (2017); Mehta and Rastegari (2022a) are introduced.
@@
Figure 1: Left: Latency comparison with Transformer Vaswani (2017), MixFormer Chen et al. (2022a), ACmix Pan et al. (2022b), Axial attention Ho et al. (2019) and local attention Luong (2015). It is measured with a single module of channel dimension 64 on a Qualcomm Snapdragon 865 processor. Right: The mIoU versus latency on the ADE20K val set. MV2 means MobileNetV2 Sandler et al. (2018). MV3-L means MobileNetV3-Large Howard et al. (2019). MV3-Lr denotes MobileNetV3-Large-reduce Howard et al. (2019). The latency is measured on a single Qualcomm Snapdragon 865, and only an ARM CPU core is used for speed testing. No other means of acceleration, e.g., GPU or quantification, is used. For figure Right, the input size is 512×512. SeaFormer achieves superior trade-off between mIoU and latency.
@@
However, these advances are still insufficient to satisfy the design requirements and constraints for mobile devices due to the high latency on the high-resolution inputs (see Figure 1). Recently there is a surge of interest in building a Transformer-based semantic segmentation. In order to reduce the computation cost at high resolution, TopFormer Zhang et al. (2022b) dedicates to applying the global attention at a $1/64$ scale of the original input, which definitely harms the segmentation performance. To solve the dilemma of high-resolution computation for pixel-wise segmentation task and low latency requirements on the mobile device in a performance-harmless way, we propose a family of mobile-friendly Transformer-based semantic segmentation models, dubbed squeeze-enhanced Axial Transformer (SeaFormer), which reduces the computational complexity of axial attention from $\mathcal{O}((H+W)HW)$ to $\mathcal{O}(HW)$, to achieve superior accuracy-efficiency trade-off on mobile devices and fill the vacancy of mobile-friendly efficient Transformer.
@@
The core building block squeeze-enhanced Axial attention (SEA attention) seeks to squeeze (pool) the input feature maps along the horizontal/vertical axis into a compact column/row and computes self-attention. We concatenate query, keys and values to compensate the detail information sacrificed during squeeze and then feed it into a depth-wise convolution layer to enhance local details. The SEA attention module is a novel enhancement designed to reduce computational complexity, making it suitable for mobile devices without sacrificing performance.
@@
Coupled with a light segmentation head and several lightweight fusion blocks, our design (see Figure 2) with the proposed SeaFormer layer in the small-scale feature is capable of conducting high-resolution image semantic segmentation with low latency on the mobile device. Our dual-branch network architecture incorporates these unique design elements that optimize efficiency and accuracy for mobile applications. As shown in Figure 1, the proposed SeaFormer outperforms other efficient neural networks on the ADE20K dataset with lower latency. In particular, SeaFormer-Base is superior to the lightweight CNN counterpart MobileNetV3 (41.0 vs.33.1 mIoU) with lower latency (106ms vs.126ms) on an ARM-based mobile device. Furthermore, we have proposed a multi-resolution distillation framework that further enhances the model’s efficiency and accuracy.
@@
A preliminary version of this work was presented in ICLR 2023 Wan et al. (2023). In this paper, we have further extended our conference version as follows: (i) We propose an adaptive squeeze and expand method in Squeeze-axial attention, using a learnable mask to map all tokens of query/key/values to a single token in each row and column; (ii) We present a feature upsampling based multi-resolution distillation approach, further reducing the inference latency of the proposed framework.
@@ 2.1 Combination of Transformers and convolution
Convolution is relatively efficient but not suitable to capture long-range dependencies and vision Transformer has the powerful capability with a global receptive field but lacks efficiency due to the computation of self attention. In order to make full use of both of their advantages, MobileViT Mehta and Rastegari (2022a), TopFormer Zhang et al. (2022b), LVT Yang et al. (2022a), Mobile-Former Chen et al. (2022b), EdgeViTs Pan et al. (2022a), MobileViTv2 Mehta and Rastegari (2022b), EdgeFormer Zhang et al. (2022a) and EfficientFormer Li et al. (2022) are constructed as efficient ViTs by combining convolution with Transformers. MobileViT, Mobile-Former, TopFormer and EfficientFormer are restricted by Transformer blocks and have to trade off between efficiency and performance in model design. LVT, MobileViTv2 and EdgeViTs keep the model size small at the cost of relatively high computation, which also means high latency.
@@ 2.2 Axial attention and variants
Axial attention Huang et al. (2019b); Ho et al. (2019); Wang et al. (2020a) is designed to reduce the computational complexity of original global self-attention Vaswani (2017). It computes self-attention over a single axis at a time and stacks a horizontal and a vertical axial attention module to obtain the global receptive field. Strip pooling Hou et al. (2020) and Coordinate attention Hou et al. (2021) uses a band shape pooling window to pool along either the horizontal or the vertical dimension to gather long-range context. Kronecker Attention Networks Gao et al. (2020) uses the juxtaposition of horizontal and lateral average matrices to average the input matrices and performs attention operation. These methods and other similar implementations provide performance gains partly at considerably low computational cost compared with Axial attention. However, they ignore the lack of local details brought by the pooling/average operation.
@@ 2.3 Mobile semantic segmentation
The mainstream of efficient segmentation methods are based on lightweight CNNs. DFANet Li et al. (2019) adopts a lightweight backbone to reduce computation cost and adds a feature aggregation module to refine high-level and low-level features. ICNet Zhao et al. (2018) designs an image cascade network to speed up the algorithm, while BiSeNet Yu et al. (2018, 2021) proposes two-stream paths for low-level details and high-level context information, separately. Fast-SCNN Poudel et al. (2019) shares the computational cost of the multi-branch network to yield a run-time fast segmentation CNN. TopFormer Zhang et al. (2022b) presents a new architecture with a combination of CNNs and ViT and achieves a good trade-off between accuracy and computational cost for mobile semantic segmentation. However, it is still restricted by the heavy computation load of global self-attention.
2.4 Knowledge distillation
Knowledge Distillation (KD) is a widely used technique in model compression, where a smaller model (student) is trained to mimic the behavior of a larger, well performing model (teacher). The concept was first introduced by Hinton et al. Hinton (2015), where the student model learns from the teacher’s softened output probabilities. Since then, various distillation methods have been proposed Romero et al. (2014); Park et al. (2019); Tian et al. (2019); Liu et al. (2022); He et al. (2019); Liu et al. (2019); Yang et al. (2022b); Cho and Hariharan (2019); Heo et al. (2019); Kim et al. (2018); Li et al. (2023c, 2021c); Mirzadeh et al. (2020); Yang et al. (2019); Yim et al. (2017); Zhang et al. (2018); Zhao et al. (2022); Wang et al. (2024) to improve the learning process, focusing on feature-based distillation Romero et al. (2014), relational knowledge distillation Park et al. (2019) and contrastive and adversarial approaches to distillation Tian et al. (2019). Our work builds on these great works by proposing a multi-resolution distillation method that aligns features at different feature resolutions between teacher and student, enabling a more efficient and effective knowledge transfer in dense prediction tasks, such as segmentation.
@@ 3.1 Overall architecture
Inspired by the two-branch architectures Yu et al. (2021); Poudel et al. (2019); Hong et al. (2021); Huang et al. (2021b); Chen et al. (2022b), we design a squeeze-enhanced Axial Transformer (SeaFormer) framework. As is shown in Figure 2, SeaFormer consists of these parts: shared STEM, context branch, spatial branch, fusion block and light segmentation head. For a fair comparison, we follow TopFormer Zhang et al. (2022b) to design the STEM. It consists of one regular convolution with stride of 2 followed by four MobileNet blocks where stride of the first and third block is 2. The context branch and the spatial branch share the produced feature map, which allows us to build a fast semantic segmentation model.
@@ 3.2 Squeeze-enhanced Axial attention
To improve the efficiency, there are some works Liu et al. (2021); Huang et al. (2019b); Ho et al. (2019) computing self-attention within the local region. We show two most representative efficient Transformer in Equation 23. Equation 2 is represented by window-based attention Luong (2015) successfully reducing the time complexity to $\mathcal{O}(m^{2}HW(C_{qk}+C_{v}))=\mathcal{O}(HW)$, where $\mathcal{N}_{m\times m}(o)$ means the neighbour $m\times m$ positions of $o$, but loosing global receptiveness. The Equation 3 is represented by Axial attention Ho et al. (2019), which only reduces the time complexity to $\mathcal{O}((H+W)HW(C_{qk}+C_{v}))=\mathcal{O}((HW)^{1.5})$, where $\mathcal{N}_{H\times 1}(o)$ means all the positions of the column of $o$; $\mathcal{N}_{1\times W}(o)$ means all the positions of the row of $o$.
@@ 3.2.5 Architecture details and variants
Figure 5: The overall pipeline of multi-resolution distillation based on feature up-sampling. MV2(E=4) denotes MobileNetV2 block with an expansion ratio of 4, and the default kernel size for depth-wise convolution is 5.
@@
Table 1: Architectures for semantic segmentation. [Conv, 3 ,16, 2] denotes regular convolution layer with kernel of 3, output channel of 16 and stride of 2. [MB, 3, 4, 16, 2] means MobileNetV2 Sandler et al. (2018) block with kernel of 3, expansion ratio of 4, output channel of 16 and stride of 2. [Sea, 2, 4] refers to SeaFormer layers with number of layers of 2 and heads of 4.
@@ 3.3 Multi-resolution distillation based on feature up-sampling
In dealing with dense prediction tasks, accurately extracting semantic information and spatial details from images often necessitates the input of high-resolution images. This approach undoubtedly increases the computational burden on the model, especially in resource-constrained environments such as mobile devices or real-time applications, where the demand for high computation becomes even more significant. In recent years, to alleviate this issue, some research Qi et al. (2021); Hu et al. (2022) has proposed methods utilizing multi-scale distillation techniques. In these methods, a teacher model that takes high-resolution images as input is used to guide a student model that takes low-resolution images as input to learn. This strategy allows the student model to produce reasonable results even with low-resolution inputs, thereby significantly reducing computational costs and increasing inference speed.
@@
Furthermore, the upscaled features from the student model are not only used for alignment with the teacher model’s features but are also integrated into the overall semantic segmentation framework, serving as spatial branch features in the feature fusion process. The specific method of this feature fusion and the involved modules are detailed in Section 3.1. Through this series of steps, the fused features are finally fed into a lightweight segmentation head to complete the semantic segmentation task.
@@ 3.4.2 Loss function
Table 2: Results of semantic segmentation on ADE20K val set, * indicates training batch size is 32. KD means knowledge distillation. The latency is measured on a single Qualcomm Snapdragon 865 with input size 512×512, and only an ARM CPU core is used for speed testing. MobileNetV3-Lr means MobileNetV3-Large-reduce. HRNet-W18S means HRNet-W18-Small. References: MobileNetV2 Sandler et al. (2018), MobileNetV3 Howard et al. (2019), HRNet Yuan et al. (2020), TopFormer Zhang et al. (2022b), PP-MobileSeg Tang et al. (2023), ConvMLP Li et al. (2023a), Semantic FPN Kirillov et al. (2019), EfficientNet Tan and Le (2019), DeepLabV3+ and Lite-ASPP Chen et al. (2018a), SegFormer Xie et al. (2021), ResNet He et al. (2016), ShuffleNetV2-1.5x Ma et al. (2018).
@@ 4 Experiments
We evaluate our method on semantic segmentation, image classification and object detection tasks. First, we describe implementation details and compare results with related efficient neural networks. We then conduct a series of ablation studies to validate the design of SeaFormer. Each proposed component and important hyper-parameters are examined thoroughly.
@@ 4.1.1 Dataset
We perform semantic segmentation experiments over ADE20K Zhou et al. (2017), CityScapes Cordts et al. (2016), Pascal Context Mottaghi et al. (2014) and COCO-Stuff Caesar et al. (2018). The mean of intersection over union (mIoU) is set as the evaluation metric. We convert full-precision models to TNN Contributors (2019) and measure latency on an ARM-based device with a single Qualcomm Snapdragon 865 processor.
@@ 4.1.2 Implementation details
We set ImageNet-1K Deng et al. (2009) pretrained network as the backbone, and training details of ImageNet-1K are illustrated in the last subsection. For semantic segmentation, the standard BatchNorm Ioffe and Szegedy (2015) layer is replaced by synchronized BatchNorm. Our implementation is based on public codebase mmsegmentation Contributors (2020). We follow the batch size, training iteration scheduler and data augmentation strategy of TopFormer Zhang et al. (2022b) for a fair comparison.
@@
The initial learning rate is 0.0005 for batch size 32 or 0.00025 for batch size 16. The weight decay is 0.01. A poly learning rate scheduled with factor 1.0 is adopted.
@@ 4.2 Comparison with state of the art
Table 2 shows the results of SeaFormer and previous efficient backbones on ADE20K val set. The comparison covers Params (model parameters), FLOPs (floating point operations), Latency and mIoU. As shown in Table 2, SeaFormer outperforms these efficient approaches with comparable or less FLOPs and lower latency. Compared with specially designed mobile backbone, TopFormer, which sets global self-attention as its semantics extractor, SeaFormer achieves higher segmentation accuracy with lower latency. And the performance of SeaFormer-B++ surpasses MobileNetV3 by a large margin of +8.3% mIoU with lower latency (-16%). The results demonstrate that our SeaFormer layers improve the representation capability significantly.
@@
We compare SeaFormer with the previous approaches on COCO-Stuff validation set in Table 5. From the results, it can be seen that SeaFormer-S++ is +1.2% mIoU higher (34.04 vs.32.82) than SeaFormer-S with a similar computation cost.